(cot^2 betta)/(csc betta)=csc betta-sin betta
1
Expert's answer
2012-02-23T10:11:46-0500
let betta=x for easy writing of
expressions

cot^2(x)/csc(x)=[cos^2(x)/sin^2(x)]/[1/sin(x)]=cos^2(x)/sin(x)=[1-sin^2(x)]/sin(x)=1/sin(x)-sin(x)=csc(x)-sin(x)